ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60K35, 82B44 Liemant, Alfred, Weierstraß-Institut Berlin, Germany Hopping transport of charge carriers on a disordered lattice. The parameters of the microscopic hopping dynamics are the lattice site concentration, the energy state density and the jump rate obeying a detailed balance. On a large time–space scale, within a mean field approximation a non–linear drift–diffusion equation is derived. Diffusivity and mobility are calculated from the microscopic characteristics of the hopping dynamics.